OpenStudy (anonymous):

what would be the first step to solve this problem 4-2d _____ =6 5

OpenStudy (mimi_x3):

\[\frac{4-2d}{5}=6\] Cross multiply

OpenStudy (mimi_x3):

\[4-2d = 6*5\]

OpenStudy (anonymous):

multiply both sides by 5

OpenStudy (mimi_x3):

\[4-2d = 30\] \[-2d = 30-4\]

OpenStudy (mimi_x3):

\[-2d = 26\]

OpenStudy (anonymous):

iWould Add Multiple 6 And 5 Like 6*5=30 And Now The Answer iS 30 Because 30 Divided By 5 = 6 Then Add 30+4 Which iS 34 And Find What You Need 2 Multiply 2 By Which iS 17 So D = 17 Beacuase 17*2=34-4=30 And 30 Divided By 5 = 6 Tada Lol :)